OpenStudy (lgbasallote):

how do you know if a radical is in its simplest form? i mean is there like a condition to be met?

OpenStudy (anonymous):

what's a radical?!

OpenStudy (lgbasallote):

 those things

OpenStudy (anonymous):

ohk

OpenStudy (lgbasallote):

do you know how?

OpenStudy (anonymous):

well these kinda things.....i dont think thr wud be a condition...jus trail nd error

Parth (parthkohli):

A radical is in its simplest form when you can't factorize the inside the radical thing and simplify it

OpenStudy (lgbasallote):

i see...anything else?

OpenStudy (lgbasallote):

it says in this paper that there are three conditions

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{10} = \sqrt{2 \times 5} = \sqrt{10} }\) You just can't simplify it more

Parth (parthkohli):

Are you trying to test us?lol

OpenStudy (lgbasallote):

uhmm no

Parth (parthkohli):

A fraction with radicals is simplified when the denominator is rationalized

OpenStudy (lgbasallote):

i think the last one had something to do with indeces but i cant remember it :/

Parth (parthkohli):

Lol I dunno about that but my logic applies to all

OpenStudy (lgbasallote):

or was it when there isnt a radical within the radical

Parth (parthkohli):

Can you give an example?

OpenStudy (lgbasallote):

well for example simplify \[\sqrt{10x^3y^5} \times \sqrt[3]{4x^2y}\]

Parth (parthkohli):

factor and simplify

Parth (parthkohli):

In the first one you can cancel out x and y.....

OpenStudy (lgbasallote):

well hmm i see that...what next? after putting out the root i meea

OpenStudy (lgbasallote):

mean*

Parth (parthkohli):

What did you get in the roots?

OpenStudy (lgbasallote):

\[\large xy \sqrt{10xy} \times \sqrt[3]{4x^2y}\] is that right?

OpenStudy (lgbasallote):

oops y should be squared in the xy part i think

Parth (parthkohli):

Lol yes

OpenStudy (lgbasallote):

then what?

Parth (parthkohli):

\( \color{Black}{\Rightarrow xy^2 \times \sqrt{10} \times \sqrt{x} \times \sqrt{y }......etc }\)

OpenStudy (lgbasallote):

i dont get what you're pointing out

Parth (parthkohli):

Do you know how to get these questions? for example: \( \color{Black}{\Rightarrow \sqrt{x} \times \sqrt[3]{x} }\)

OpenStudy (lgbasallote):

uhmm...x^1/6?

Parth (parthkohli):

My personal preference is to use exponents. \( \color{Black}{\Rightarrow x^{1 \over 2} \times x^{1 \over 3} }\) You add the exponents when base is same and you are multiplying

OpenStudy (lgbasallote):

oh oh yeah add exponENTS *facepalm* well what about in 10 and 4? it's not same base anymore

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{10} \times \sqrt[3]{4} }\) Some thing like this?

OpenStudy (lgbasallote):

then? how do you express it as one index??

Parth (parthkohli):

Lol here you are a free bird.....use the internet to figure out

Parth (parthkohli):

Oh I see how to do this yes

OpenStudy (lgbasallote):

i come here to learn and you're directing me to the net?!

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{10} = \sqrt{2} \times \sqrt{5} }\)

OpenStudy (lgbasallote):

i thought the purpose of this site was to teach people???

OpenStudy (lgbasallote):

i see no logic in telling people to go away and learn from the net...the purpose i came here instead of the net is because i didnt find my answer in the net right?!

Parth (parthkohli):

Hmm but I just told you

OpenStudy (lgbasallote):

i still dont see your point btw

Parth (parthkohli):

Do you know that:; \( \color{Black}{\Rightarrow \sqrt{a} \times \sqrt{b} = \sqrt{ab} }\)?

OpenStudy (lgbasallote):

yeah but how does iot apply to 10 and 4

Parth (parthkohli):

Now look: \( \color{Black}{\Rightarrow \sqrt{10} = \sqrt{2} \times \sqrt{5} }\) agree?

OpenStudy (lgbasallote):

yeah

Parth (parthkohli):

Aww snap!

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt[3]{4} = \sqrt[3]{2} \times \sqrt[3]{2} }\) Agree again?

OpenStudy (lgbasallote):

yes

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{2} \times \sqrt{5} \times\sqrt[3]{2} \times \sqrt[3]{2} }\) Convert into exponents and see :D

OpenStudy (lgbasallote):

what??

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2^{1 \over 2} \times 5^{1 \over 2} \times 2^{1 \over 3} \times 2^{1 \over 3} }\)

Parth (parthkohli):

Add the exponents of 2.

OpenStudy (lgbasallote):

10^1/2 you just went back to the old one lol..long cut

Parth (parthkohli):

So do you understand how to do it?

OpenStudy (lgbasallote):

nope..what's after 10^1/2

Parth (parthkohli):

Lol I told you that: \( \color{Black}{\Rightarrow \sqrt{10} = \sqrt{2} \times \sqrt{5} = 2^{1 \over 2} \times 5^{1 \over 2} }\)

OpenStudy (lgbasallote):

you're circling me around you know

Parth (parthkohli):

uh-oh....you're getting circled. Twid me

OpenStudy (lgbasallote):

okay...